# Causality Pursuit from Heterogeneous Environments via Neural Adversarial Invariance Learning

### v1

@@ A.5 Discussion on the Nondegenerate Intervention Condition
We suppose the following condition similar to 4.8 holds in the constructed graph.
Condition A.1.
@@
We are ready to give a complete statement of Proposition 4.4.
Proposition A.3 (Formal Statement of Proposition 4.4).
Under the setting of Theorem 4.3, for a new environment $t$ with SCM $M^{(t)}=\{\mathcal{S}^{(t)},\nu\}$ satisfying $f_{j}^{(t)}\equiv f_{j}^{(0)}$ for any $j\in[d+1]\setminus I$, i.e., only $X_{I}$ is intervened, we also have $\mathbb{E}[Y^{(t)}|X_{S_{\star}}^{(t)}]\equiv\mathbb{E}[Y^{(0)}|X_{S_{\star}}^{(0)}]$. Suppose further that A.1 holds for the constructed SCM $\bar{M}$. Then $S_{\star}$ is the unique largest set whose conditional expectation is transferable, i.e., for any $S\subseteq[d]$ such that $\mathbb{E}[Y^{(t)}|X_{S_{\star}\cup S}^{(t)}]\neq\mathbb{E}[Y^{(t)}|X_{S_{\star}}^{(t)}]$, one has $\mathbb{E}[Y^{(t)}|X_{S}^{(t)}]\neq\mathbb{E}[Y^{(0)}|X_{S}^{(0)}]$.
A.7 Technical Related Works
Adversarial Estimation. Adversarial estimation is introduced in Goodfellow et al., (2014) for generative modeling. There is some literature on formalizing the adversarial estimation from a statistical perspective and providing statistical analysis for this family of generative modeling methods (Liang,, 2021; Singh et al.,, 2018; Biau et al.,, 2020). It is also used as a tool for estimating conditional moment models, for example, instrumental variable regression (Dikkala et al.,, 2020), estimating the (implicit) influence function (Chernozhukov et al.,, 2020; Hirshberg & Wager,, 2021).
@@ B.1 Main Result for the General FAIR Least Squares Estimator
This section is designed to offer a unified main result characterizing when the FAIR least squares estimator can identify the target regression function together with a non-asymptotic $L_{2}$ error bound for general $(\mathcal{G},\mathcal{F})$. To impose statistical complexity on the function classes we used, we introduce the definition of localized population Rademacher complexity, described as follows.
@@
Condition B.1 (Function Class).
@@
Note that when $-\mathcal{G}=\mathcal{G}$, $R_{n,\mu^{(e)}}(\delta;\partial\mathcal{G})=R_{n,\mu^{(e)}}(\delta;\mathcal{G})$. The above three assumptions 4.1, 4.3, B.1 are standard in the theoretical analysis of regression. Now, we introduce the specific assumption in our multi-environment regression setting.
Condition B.2 (Invariance and Identification).
@@
The second condition “heterogeneity” is for identification and is fundamental to derive the population-level strong convexity with respect to $g^{\star}$. The two quantities in (B.1) are general forms of the bias mean and the bias variance, respectively. We refer to $\mathsf{b}_{\mathcal{G}}(S)$ as the bias mean because $\mathsf{b}_{\mathcal{G}}(S)$ is the precise bias of the estimator that regress $Y$ on $X_{S}$ when $S^{\star}\subseteq S$ using all the data. This can be formally presented in the following proposition, which asserts that in the absence of our proposed regularizer, a vanilla least squares estimator will not consistently estimate $g^{\star}$.
Proposition B.1 (Inconsistency of Least Squares Estimator).
Let $S$ be an index set such that $S^{\star}\subseteq S\subseteq[d]$. Assume 4.1, 4.3, B.1B.2 hold, and $\mathsf{b}_{\mathcal{G}}(S)>0$. Suppose further that $U\delta_{n,\log n}+\inf_{g\in\mathcal{G}_{S}}\|g-\Pi_{\overline{\mathcal{G}_{S}}}(\bar{m}^{(S)})\|_{2}=o(1)$, where $U$ and $\delta_{n,t}$ are two constants defined in Theorem B.2 below. Then the estimator $\widehat{g}_{\mathtt{R}}$ that minimizes (3.6) in $\mathcal{G}_{S}$ satisfies, for large enough $n$,
@@
On the other hand, our proposed FAIR estimator will not converge to the biased solution under the condition “heterogeneity”. The condition “heterogeneity” is an abstraction of the “identification” condition in previous subsections, for example, 4.5 for FAIR-NN.
@@
Now we are ready to state the main result regarding the statistical rate of convergence of our estimator $\widehat{g}$ to $g^{\star}$, that is,
@@
Theorem B.2 (Main Result for the FAIR Estimator with $\ell_{2}$ Loss).
Assume 4.1, 4.3, and B.1B.2 hold. Define the critical threshold
@@
and the stochastic error as $\delta_{n,t}=\delta_{n}+\{(\log(nB|\mathcal{E}|)+t+1)/n\}^{1/2}$, where $\delta_{n}$ is the quantity in B.1. Let $U=B(B+\sigma_{y}\sqrt{\log(n|\mathcal{E}|)})$, then
@@
Theorem B.2 generalizes Theorem 4.4 in Fan et al., (2023) to a broad spectrum of $(\mathcal{G},\mathcal{F})$ configurations. After specifying the function class $(\mathcal{G},\mathcal{F})$, one can further derive the corresponding identification condition by calculating $(\mathsf{b}_{\mathcal{G}}(S),\bar{\mathsf{d}}_{\mathcal{G},\mathcal{F}}(S))$ and establish a high probability bound on the $L_{2}$ error by substituting approximation errors $(\delta_{\mathtt{a},\mathcal{G}},\delta_{\mathtt{a},\mathcal{F},\mathcal{G}}(S),\delta_{\mathtt{a},\mathcal{F},\mathcal{G}}^{\star})$ and stochastic error $\delta_{n}$ for the function class $(\mathcal{G},\mathcal{F})$. In particular, when $\mathcal{G}$ and $\mathcal{F}$ are restricted to the linear function class, they not only match but also significantly improve the result in Fan et al., (2023); see Section B.6. All the results in Table 2 are direct corollaries of our abstract result Theorem B.2.
It is required that $\gamma$ should be greater than a constant-level critical threshold $8\gamma^{\star}$ for consistent estimation of $g^{\star}$. Theorem B.2 further establishes a crude instant-dependent and oracle-type error bound (B.2) that holds for arbitrary $n\geq 2$ and scales linearly with $\gamma$. Furthermore, when the stochastic error and approximation errors all go to 00 as $n$ increases and $n$ is large enough such that (B.3) holds, we have (B.4), which improves the $L_{2}$ error bound (B.2) in two aspects – the error bound is no longer dependent on either $\gamma$ or other $m^{(e,S)}$ with $S\neq S^{\star}$. The quantities in the RHS of (B.3) can be interpreted as the smaller of (1) the signal of true important variables and (2) the signal of heterogeneity. When one of these signals is weak, one can expect to demand more data to differentiate whether it is signal or noise.
One important ingredient in the FAIR estimator is the choice of regularization hyper-parameter $\gamma$ that promotes the invariance. Theorem B.2 offers some insights on choosing $\gamma$. Firstly, $\gamma\geq C\gamma^{\star}$ is required such that it will correctly identify $g^{\star}$ from a population-level perspective. Second, it will influence the $L_{2}$ error rate when $n$ is not large enough such that (B.3) does not hold. Furthermore, the final $L_{2}$ error rate (B.4) when $n$ is large enough is independent of $\gamma$. This indicates that the estimator’s performance is somewhat not very sensitive to the choice of hyper-parameter $\gamma$. In this case, one can adopt a slightly conservative large $\gamma$ to meet the population condition $\gamma\geq C\gamma^{\star}$.
B.2 Extension to the General Risk Loss $\ell$ under the Nonparametric Setting
Condition B.3 (Risk Loss).
@@ B.2 Extension to the General Risk Loss $\ell$ under the Nonparametric Setting
The assumptions on risk loss in B.3 is standard: (1) ensures that $\ell$ is well-defined on optimal solutions and linear combination of them, (2) requires that the population-level global minima is conditional mean, (3) guarantees that the loss function is strongly convex and smooth in the domain, and satisfies $|\ell(y,v)-\ell(y,v^{\prime})|\leq\zeta|y-\widetilde{v}||v-v^{\prime}|$ for some universal constant $\zeta$, which slightly relaxes the Lipschitz condition in Farrell et al., (2021) and Foster & Syrgkanis, (2019).
@@
Condition B.4 (Invariance and Identification for General Risk Loss).
@@
Theorem B.3 (Main Result for the FAIR Estimator with General Risk Loss).
Assume 4.1,4.3, B.1, and B.3B.4 hold. Define the critical threshold
@@
and the stochastic error as $\delta_{n,t}=\delta_{n}+\{(\log(nB|\mathcal{E}|)+t+1)/n\}^{1/2}$, where $\delta_{n}$ is the quantity in B.1. Let $U=B(B+\sigma_{y}\sqrt{\log(n|\mathcal{E}|)})$, then
@@
B.3 Key Ideas and Proof Sketch of Theorem B.2
@@
Theorem B.4.
Assume B.2 hold, $\ell(y,v)=\frac{1}{2}(y-v)^{2}$. Let $\delta\in(0,1)$ be arbitrary. Then the following holds, for any $\gamma\geq 4\delta^{-1}\gamma^{\star}$,
@@
Proposition B.5 (Instance-dependent error bounds for pooled risk).
Suppose 4.1,4.3, B.1 hold. There exists some universal constant $C$ such that for any $\eta>0$ and $t>0$, the following event
@@
Proposition B.6 (Instance-dependent error bounds for regularizer).
Suppose 4.1, 4.3, B.1 hold. There exists some universal constant $C$ such that for any $t>0$, the following event
@@
We first utilize Proposition B.6 in a way that $g$ and $\widetilde{g}$ are the same. In this case, the optimization problem of $\max$$-\mathcal{F}$ in one single environment $e\in\mathcal{E}$ for fixed $g\in\mathcal{G}$ is similar to least squares regression that fits the target regression function
@@
We formally present the above intuition in the following instance-dependent error bound in Proposition B.7 in a way that the optimization gap term is maintained in the error bound.
Proposition B.7 (Instance-dependent characterization of approximately optimal discriminator).
Let $20<\eta<1/$ 2 be arbitrary, under the event defined in Proposition B.6, the following holds,
@@
where $C$ is the universal constant defined in Proposition B.6. Averaging over all the $e\in\mathcal{E}$, we obtain
@@
Now we are ready to prove Theorem B.2.
@@
    In the first step, we establish a variable selection property claim that when the Eq. B.3 holds, and the events defined in Proposition B.5 and B.6 occurs, then $\widehat{S}$ satisfies
@@
    using proof by contradiction that any $g$ such that such that the above constrain is violated in $S_{g}$, will not be the approximate solution of the minimax optimization $\inf_{g}\sup_{f^{\mathcal{E}}}\widehat{\mathsf{Q}}_{\gamma}(g,f^{\mathcal{E}})$. This can be summarized as the following Proposition B.8.
@@
    In the second step, we proceed conditioned on the above claim and derive a sharp $L_{2}$ error bound. To derive a sharp error bound, we combine (1) the approximate strong convexity with respect to $g^{\star}$, i.e., Theorem B.4, (2) the instance-dependent error bound for $\mathsf{J}$ and $\mathsf{R}$, i.e., Proposition B.5 and B.6, and (3) the key fact that, if the claim in step 1 holds, then
@@
Proposition B.8.
Under the event defined in Proposition B.6 and B.5, we have the event
@@
B.4 Applications of Theorem B.2 and Connection to the Predecessors

### v2

@@ A.6 The Complete Statement of Proposition 2
We suppose the following condition similar to 5 holds in the constructed graph.
Condition 6.
@@
We are ready to give a complete statement of Proposition 2.
Proposition 5 (Formal Statement of Proposition 2).
Under the setting of Theorem 2, for a new environment $t$ with SCM $M^{(t)}=\{\mathcal{S}^{(t)},\nu\}$ satisfying $f_{j}^{(t)}\equiv f_{j}^{(0)}$ for any $j\in[d+1]\setminus I$, i.e., only $X_{I}$ is intervened, we also have $\mathbb{E}[Y^{(t)}|X_{S_{\star}}^{(t)}]\equiv\mathbb{E}[Y^{(0)}|X_{S_{\star}}^{(0)}]$. Suppose further that 6 holds for the constructed SCM $\bar{M}$. Then $S_{\star}$ is the unique largest set whose conditional expectation is transferable, i.e., for any $S\subseteq[d]$ such that $\mathbb{E}[Y^{(t)}|X_{S_{\star}\cup S}^{(t)}]\neq\mathbb{E}[Y^{(t)}|X_{S_{\star}}^{(t)}]$, one has $\mathbb{E}[Y^{(t)}|X_{S}^{(t)}]\neq\mathbb{E}[Y^{(0)}|X_{S}^{(0)}]$.
@@ B.1 Main Result for the General FAIR Least Squares Estimator
This section is designed to offer a unified main result characterizing when the FAIR least squares estimator can identify the target regression function together with a non-asymptotic $L_{2}$ error bound for general $(\mathcal{G},\mathcal{F})$. We first introduce some standard regularity conditions.
Condition 7 (Data Generating Process).
Condition 8 (Sub-Gaussian Response).
To impose statistical complexity on the function classes we used, we introduce the definition of localized population Rademacher complexity, described as follows.
@@
Condition 9 (Function Class).
@@
Note that when $-\mathcal{G}=\mathcal{G}$, $R_{n,\mu^{(e)}}(\delta;\partial\mathcal{G})=R_{n,\mu^{(e)}}(\delta;\mathcal{G})$. The above three assumptions 7, 8, 9 are standard in the theoretical analysis of regression. Recall the definition of $m^{(e,S)}$ and $\bar{m}^{(S)}$ in Section 2.1, now we introduce the specific assumption in our multi-environment regression setting.
Condition 10 (Invariance and Identification).
@@
The second condition “heterogeneity” is for identification and is fundamental to derive the population-level strong convexity with respect to $g^{\star}$. The two quantities in (B.1) are general forms of the bias mean and the bias variance, respectively. We refer to $\mathsf{b}_{\mathcal{G}}(S)$ as the bias mean because $\mathsf{b}_{\mathcal{G}}(S)$ is the precise bias of the estimator that regress $Y$ on $X_{S}$ when $S^{\star}\subseteq S$ using all the data. This can be formally presented in the following proposition, which asserts that in the absence of our proposed regularizer, a vanilla least squares estimator will not consistently estimate $g^{\star}$, and the discrepancy $\|\widehat{g}-g^{\star}\|_{2}^{2}$ is approximately equal to $\mathsf{b}(S)$ when $n$ is large.
Proposition 6 (Inconsistency of Least Squares Estimator).
Let $S$ be an index set such that $S^{\star}\subseteq S\subseteq[d]$. Assume 7, 8, 910 hold, and $\mathsf{b}_{\mathcal{G}}(S)>0$. Suppose further that $U\delta_{n,\log n}+\inf_{g\in\mathcal{G}_{S}}\|g-\Pi_{\overline{\mathcal{G}_{S}}}(\bar{m}^{(S)})\|_{2}=o(1)$, where $U$ and $\delta_{n,t}$ are two constants defined in Theorem 4 below. Then the estimator $\widehat{g}_{\mathtt{R}}$ that minimizes (4.6) in $\mathcal{G}_{S}$ satisfies, for large enough $n$,
@@
On the other hand, our proposed FAIR estimator will not converge to the biased solution under the condition “heterogeneity”. The condition “heterogeneity” is an abstraction of the “identification” condition in previous subsections, for example, 2 for FAIR-NN.
@@
$\sup_{f^{\mathcal{E}}\in\{\mathcal{F}_{S_{\widehat{g}}}\}^{|\mathcal{E}|}}\widehat{\mathsf{Q}}_{\gamma}(\widehat{g},f^{\mathcal{E}})-(\gamma+1)\delta_{\mathtt{opt}}^{2}\leq\widehat{\mathsf{Q}}_{\gamma}(\widehat{g},\widehat{f}^{\mathcal{E}})\leq\inf_{g\in\mathcal{G}}\sup_{f^{\mathcal{E}}\in\{\mathcal{F}_{S_{g}}\}^{|\mathcal{E}|}}\widehat{\mathsf{Q}}_{\gamma}(g,f^{\mathcal{E}})+(1+\gamma)\delta_{\mathtt{opt}}^{2}$ (B.2)
with some optimization error $\delta_{\mathtt{opt}}^{2}>0$, here $\gamma$ in $(1+\gamma)$ is the same as that in $\widehat{\mathsf{Q}}_{\gamma}$. Now we are ready to state the main result regarding the statistical rate of convergence of our estimator $\widehat{g}$ to $g^{\star}$, that is,
@@
Theorem 4 (Main Result for the FAIR Estimator with $\ell_{2}$ Loss).
Assume Conditions 710 hold. Define the critical threshold
@@
and the stochastic error as $\delta_{n,t}=\delta_{n}+\{(\log(nB|\mathcal{E}|)+t+1)/n\}^{1/2}$, where $\delta_{n}$ is the quantity in 9. Let $U=B(B+\sigma_{y}\sqrt{\log(n|\mathcal{E}|)})$, then
@@
Theorem 4 generalizes Theorem 4.4 in Fan et al., (2023) to a broad spectrum of $(\mathcal{G},\mathcal{F})$ configurations. After specifying the function class $(\mathcal{G},\mathcal{F})$, one can further derive the corresponding identification condition by calculating $(\mathsf{b}_{\mathcal{G}}(S),\bar{\mathsf{d}}_{\mathcal{G},\mathcal{F}}(S))$ and establish a high probability bound on the $L_{2}$ error by substituting approximation errors $(\delta_{\mathtt{a},\mathcal{G}},\delta_{\mathtt{a},\mathcal{F},\mathcal{G}}(S),\delta_{\mathtt{a},\mathcal{F},\mathcal{G}}^{\star})$ and stochastic error $\delta_{n}$ for the function class $(\mathcal{G},\mathcal{F})$. In particular, when $\mathcal{G}$ and $\mathcal{F}$ are restricted to the linear function class, they not only match but also significantly improve the result in Fan et al., (2023); see Section B.6. All the results in Table 2 are direct corollaries of our abstract result Theorem 4.
It is required that $\gamma$ should be greater than a constant-level critical threshold $8\gamma^{\star}$ for consistent estimation of $g^{\star}$. Theorem 4 further establishes a crude instant-dependent and oracle-type error bound (B.3) that holds for arbitrary $n\geq 2$ and scales linearly with $\gamma$. Furthermore, when the stochastic error and approximation errors all go to 00 as $n$ increases and $n$ is large enough such that (B.4) holds, we have (B.5), which improves the $L_{2}$ error bound (B.3) in two aspects – the error bound is no longer dependent on either $\gamma$ or other $m^{(e,S)}$ with $S\neq S^{\star}$. The quantities in the RHS of (B.4) can be interpreted as the smaller of (1) the signal of true important variables and (2) the signal of heterogeneity. When one of these signals is weak, one can expect to demand more data to differentiate whether it is signal or noise.
One important ingredient in the FAIR estimator is the choice of regularization hyper-parameter $\gamma$ that promotes the invariance. Theorem 4 offers some insights on choosing $\gamma$. Firstly, $\gamma\geq C\gamma^{\star}$ is required such that it will correctly identify $g^{\star}$ from a population-level perspective. Second, it will influence the $L_{2}$ error rate when $n$ is not large enough such that (B.4) does not hold. Furthermore, the final $L_{2}$ error rate (B.5) when $n$ is large enough is independent of $\gamma$. This indicates that the estimator’s performance is somewhat not very sensitive to the choice of hyper-parameter $\gamma$. In this case, one can adopt a slightly conservative large $\gamma$ to meet the population condition $\gamma\geq C\gamma^{\star}$.
B.2 Extension to the General Risk Loss under the Nonparametric Setting
Condition 11 (Risk Loss).
@@ B.2 Extension to the General Risk Loss under the Nonparametric Setting
The assumptions on risk loss in 11 is standard: (1) ensures that $\ell$ is well-defined on optimal solutions and linear combination of them, (2) requires that the population-level global minima is conditional mean, (3) guarantees that the loss function is strongly convex and smooth in the domain, and satisfies $|\ell(y,v)-\ell(y,v^{\prime})|\leq\zeta|y-\widetilde{v}||v-v^{\prime}|$ for some universal constant $\zeta$, which slightly relaxes the Lipschitz condition in Farrell et al., (2021) and Foster & Syrgkanis, (2019).
@@
Condition 12 (Invariance and Identification for General Risk Loss).
@@
Theorem 5 (Main Result for the FAIR Estimator with General Risk Loss).
Assume 7,8,9, and 1112 hold. Define the critical threshold
@@
and the stochastic error as $\delta_{n,t}=\delta_{n}+\{(\log(nB|\mathcal{E}|)+t+1)/n\}^{1/2}$, where $\delta_{n}$ is the quantity in 9. Let $U=B(B+\sigma_{y}\sqrt{\log(n|\mathcal{E}|)})$, then
@@
B.3 Key Ideas and Proof Sketch of Theorem 4
@@
Theorem 6.
Assume 10 hold, $\ell(y,v)=\frac{1}{2}(y-v)^{2}$. Let $\delta\in(0,1)$ be arbitrary. Then the following holds, for any $\gamma\geq 4\delta^{-1}\gamma^{\star}$,
@@
Proposition 7 (Instance-dependent error bounds for pooled risk).
Suppose 7,8, 9 hold. There exists some universal constant $C$ such that for any $\eta>0$ and $t>0$, the following event
@@
Proposition 8 (Instance-dependent error bounds for regularizer).
Suppose 7, 8, 9 hold. There exists some universal constant $C$ such that for any $t>0$, the following event
@@
We first utilize Proposition 8 in a way that $g$ and $\widetilde{g}$ are the same. In this case, the optimization problem of $\max$$-\mathcal{F}$ in one single environment $e\in\mathcal{E}$ for fixed $g\in\mathcal{G}$ is similar to least squares regression that fits the target regression function
@@
We formally present the above intuition in the following instance-dependent error bound in Proposition 9 in a way that the optimization gap term is maintained in the error bound.
Proposition 9 (Instance-dependent characterization of approximately optimal discriminator).
Let $20<\eta<1/$ 2 be arbitrary, under the event defined in Proposition 8, the following holds,
@@
where $C$ is the universal constant defined in Proposition 8. Averaging over all the $e\in\mathcal{E}$, we obtain
@@
Now we are ready to prove Theorem 4.
@@
    In the first step, we establish a variable selection property claim that when the Eq. B.4 holds, and the events defined in Proposition 7 and 8 occurs, then $\widehat{S}$ satisfies
@@
    using proof by contradiction that any $g$ such that such that the above constrain is violated in $S_{g}$, will not be the approximate solution of the minimax optimization $\inf_{g}\sup_{f^{\mathcal{E}}}\widehat{\mathsf{Q}}_{\gamma}(g,f^{\mathcal{E}})$. This can be summarized as the following Proposition 10.
@@
    In the second step, we proceed conditioned on the above claim and derive a sharp $L_{2}$ error bound. To derive a sharp error bound, we combine (1) the approximate strong convexity with respect to $g^{\star}$, i.e., Theorem 6, (2) the instance-dependent error bound for $\mathsf{J}$ and $\mathsf{R}$, i.e., Proposition 7 and 8, and (3) the key fact that, if the claim in step 1 holds, then
@@
Proposition 10.
Under the event defined in Proposition 8 and 7, we have the event
@@
B.4 Applications of Theorem 4 and Connection to the Predecessors